sinx+cosx+cosecx+secx+tanx+cotx=7 and sin2x = a-b(sqrt7), then
 
A) a=8                  B) a=22 
C)b=8                        D)b=2

Grade:12

1 Answers

Srilekha
5 years ago
This expression equals to(Sin^2 x+1)/(sin x) +(cos^2 x+1)/(cos x)+tan^2 x+1)/tan x=((sin x+cos x)(sinx cosx+1)÷(sin x cos x))+ 2/sin 2x= (1+sin 2x)(1+(sin 2x)/2)÷2/sin 2xOn substituting sin 2x=a-broot 7; and substitute the values of a and b, we get the correct answer